# Trace: Contrastive learning for multi-trial time-series data in neuroscience

### v1

@@ Abstract
Modern neural recording techniques such as two-photon imaging allow to acquire vast time-series datasets with responses of hundreds or thousands of neurons. Contrastive learning is a powerful self-supervised framework for learning representations of complex datasets. Existing applications for neural time series rely on generic data augmentations and do not exploit the multi-trial data structure inherent in many neural datasets. Here we present Trace, a new contrastive learning framework that averages across different subsets of trials to generate positive pairs. Trace allows to directly learn a two-dimensional embedding, combining ideas from contrastive learning and neighbor embeddings. We show that Trace outperforms other methods, resolving fine response differences in simulated data. Further, using in vivo recordings, we show that the representations learned by Trace capture both biologically relevant continuous variation, cell-type-related cluster structure, and can assist data quality control.
@@ 1 Introduction
With advances in recording techniques, datasets in neuroscience have grown in size and complexity [7, 36]. For example, two-photon imaging and Neuropixel probes have made it possible to record responses of tens of thousands of neurons from multiple cortical areas under comparable conditions [34, 12]. To summarize and visually explore such noisy, high-dimensional data, it is invaluable to represent it in two dimensions to identify functionally similar groups of neurons [44].
@@
A prominent paradigm for learning informative representations of data is contrastive learning [16, 27]. Here, representations are created by contrasting similar samples (referred to as “positive pairs”) against dissimilar ones (“negative pairs”), ensuring that similar samples are grouped together, while dissimilar ones are separated. Contrastive learning has been popular for image data for some years [8], but has only recently seen first applications for neuroscience time-series data [33, 40]. For example, the contrastive method CEED [40] learns representations of extracellular action potential wave forms using generic data augmentations such as amplitude jitter. However, most contrastive learning methods, including CEED, do not directly embed into two dimensions necessary for visualization.
@@
Here, we exploit a common feature of neuroscience experiments – multiple recorded responses to identical stimuli – to develop a new contrastive learning framework tailored for multi-trial time series in neuroscience. This framework directly learns a two-dimensional (2D) representation of the data. Building on $t$-SimCNE [6], we present TraceTime series Representation Analysis through Contrastive Embeddings. Instead of using hand-tuned or generic data augmentations, Trace creates positive pairs by averaging across different subsets of trials, better capturing the structure of time-series variation in neural responses (Fig. 1).
@@
Figure 1: Trace embeds multi-trial time series using subset means for positive-pair generation. Left: Typical experimental structure in neuroscience: multiple trials of neural activity in response to repeats of an identical stimulus (here, full-field light intensity modulations). For Trace, positive pairs are generated using means of subsets of trials. Middle: Positive pairs are fed through an MLP and a fully-connected projection head to get representations $z_{i}$ and $z^{\prime}_{i}$. The loss function pushes $z_{i}$ and $z^{\prime}_{i}$ together and maximizes their similarity. Right: Final embedding of large-scale neuroscience dataset recorded in superior colliculus. Color-coded according to their functional group.
@@
Using a synthetic dataset, we show that Trace indeed automatically identifies informative parts of a time series, where other methods are “drowning in noise”. We then investigate the performance of Trace on a two-photon-imaging dataset of over 70,000 individual retinal ganglion cell boutons recorded in the superior colliculus of awake, head-fixed mice. Trace produces visualizations that capture important biological properties of the recorded cells better than other visualization techniques.
@@
Our code is available at
https://osf.io/x7qsm/?view_only=e5fd82164c8a4a30807cdba400d80c28.
@@ 2 Related work
Towards this goal, two of the most prominent non-linear visualization methods are $t$-SNE [39] and UMAP [26]. While mostly used for single-cell transcriptomics data, they have also been employed for wave-shape classification for electrophysiological data [21]. $t$-SNE and UMAP learn a non-parametric 2D embedding, guided by pairs of nearest neighbors in data space. However, nearest neighbors in the high-dimensional space can be a poor proxy for semantic similarity due to the curse of dimensionality [17]. Alternatively, self-supervised learning approaches train encoder networks to produce high-dimensional representations. For visual exploration, these need to be further reduced, e.g. using PCA [30]. Instead of pairs of nearest neighbors, these methods rely on data augmentations, better capturing the semantic similarities of samples. The type of augmentation depends on the data modality, e.g. randomly resized crops and flips for images [8] or sampling consecutive time steps in speech recordings [27]. Notably, $t$-SimCNE [6] uses data augmentations to create positive pairs of images but embeds into 2D to visualize the data.
@@
Contrastive learning has been applied to medical time-series data, in particular electrocorticogram (ECoG) data. In addition to general time-series augmentations like scaling, blurring [32], and jittering [15], frequency- [46], cutout- [9], and permutation-based augmentations [32] have been used, which are not applicable to multi-trial non-periodic, stimulus-response data. Some of these works [19, 13] also create positive pairs based on patient identity. Mix-up [45] creates augmentations by blending different samples [4], which was also applied to ECoG data [9, 41]. Patient-based positive pairs or mix-up are similar in spirit to our approach. An important difference is that they only mix or pair two single trial samples and not a larger subset of trials, as we do. Our approach balances positive pair variability and similarity to inference-time inputs.
@@
In neuroscience, there has been comparatively little work on contrastive learning for time-series representations with the goal of identifying discrete cell types (i.e. clusters). Peterson et al. [31] use a self-supervised model to generate pseudo labels and enrich the collected data this way. Cho et al. [10] suggest a time-warping loss, which is not applicable to our experiments, which are aligned in time by construction. Related to our approach, the method CEED has been used for spike sorting and cell-type classification [40]. CEED generates representations using generic time-series as well as task-specific augmentations such as spike collision and channel subset selection. As is common for contrastive self-supervised learning, CEED does not embed into two dimensions directly, but relies on an additional dimensionality reduction step for a 2D visualization, unlike our method. Moreover, Trace captures the local noise structure better than CEED’s generic augmentations (Fig. 2).
@@
For learning representations of neural population dynamics, methods like LFADS [29], SwapVAE [24], or the contrastive approaches MYOW [3], CEBRA [33], SinkDivLM [38], and Neuroformer [2] produce embeddings in which the temporal evolution of the population activity and behavioral modulation of neural activity can be explored. Here, each embedding point corresponds to a time point, not a single-neuron response like in Trace. Thus, they tackle a fundamentally different problem and are not applicable in our setting (see point (2) in the beginning of this section).
@@
Similarly, many general-purpose contrastive time-series models, like TNC [37], also produce time-point embeddings and are thus not applicable in our setting. Others, such as TS2Vec [43] contrast instances of time series, but learn high-dimensional embeddings, not geared towards visualization. We find that Trace outperforms TS2Vec even when adapting the latter to a visualization setting.
@@
NEMO [42] and PhysMAP [22] contrast different electorphysiological modalities, while Trace works by contrasting multiple recordings of the same modality.
@@ 3 Trace: contrastive learning for multi-trial time-series data in neuroscience
Our goal is to create informative 2D embeddings of multi-trial time-series data from neuroscience experiments, supporting data exploration, cell-type discovery, and other clustering tasks. To do so, we developed Trace, a method that makes use of the trial-based nature of many neuroscience experiments to create positive pairs by averaging over subsets of trials. It builds on the self-supervised method $t$-SimCNE [6], which is based on the SimCLR framework [8] and directly learns a 2D embedding. We briefly review contrastive learning and $t$-SimCNE before describing our approach. Finally, we will briefly describe the contrastive learning frameworks CEED and TS2Vec as our closest competitors using generic data augmentations for time series [40].
Figure 2: Synthetic dataset shows ability of Trace to identify cell types in noisy data. (a) Simulation of two classes of neural responses with baseline (25 s) and type-specific response (5 s) periods. In the baseline period, both classes show 10 Hz activity, while in the type-specific response period they fire at 11 Hz and 9 Hz, respectively. Responses are corrupted by variable amounts of noise (Sec. 4.1). (b) Discriminability [14] of the two neuron classes per bin was measured using $D=\frac{|\mu_{1}-\mu_{2}|}{0.5(\sigma_{1}+\sigma_{2})}$ (mean across neurons shown). (c) Mean neural responses across 10 trials for two example neurons from the two types. (d) Distributions of baseline (dashed) and response periods (solid) for the two neuronal responses in (c). (e – i) Embeddings of 10k simulated neurons for baseline activity noise SD={}={}=38 with 10 trials each. (j) Discriminability for simulated responses with increasing amounts of noise during the neural baseline activity. Error bars indicate 95% confidence intervals.
@@ 3.1 Contrastive self-supervised visualizations
Given a dataset $\{x_{1},\dots,x_{n}\}\subset X$ in a data space $X$, contrastive self-supervised learning trains the parameters $\theta$ of a neural network $f_{\theta}:X\to Z$ to obtain representations $z_{i}=f_{\theta}(x_{i})$ in embedding space $Z$. It learns salient features by making the embedding invariant to known data similarities encoded in positive pairs. Similarity in the embedding space is measured by a similarity function $q_{ij}=q(z_{i},z_{j})$, Equation 2. Positive pairs $ii^{\prime}$ should have high similarity $q_{ii^{\prime}}$, i.e. $z_{i}$ and $z_{i^{\prime}}$ should be close. This is typically achieved with the InfoNCE loss, which reads for positive pair $ii^{\prime}$
$\mathcal{L}_{ii^{\prime}}=-\log\frac{q_{ii^{\prime}}}{q_{ii^{\prime}}+\sum_{\alpha\neq i,i^{\prime}}q_{i\alpha}}.$ (1)
Here, $\alpha$ runs over all elements in the training batch other than those of the positive pair. The pairs $i\alpha$ are called negative pairs and their similarity is decreased, i.e. $z_{i}$ and $z_{\alpha}$ pushed apart, to prevent a collapsed solution, mapping the entire dataset to a single point.
Usually, the embedding space is constrained to a high-dimensional hypersphere $Z=S^{d-1}$ and the similarity function is based on the cosine similarity:
$q_{ij}=\exp\left(\frac{z_{i}^{\top}z_{j}}{\|z_{i}\|\|z_{j}\|\tau}\right),$ (2)
@@
$t$-SimCNE [6] learns a 2D visualization of image datasets with a parametric encoder and data augmentations as similarity source. The key ingredient is the $t$-SNE-inspired Cauchy kernel
$q_{ij}=(1+\|z_{i}-z_{j}\|^{2})^{-1}$ (3)
in the 2D embedding space $Z=\mathbb{R}^{2}$. We adapt $t$-SimCNE to multi-trial time-series data from neuroscience by implementing time-series-specific data augmentations, replacing the original image-based transformations (see below). We adopt a simpler (one-stage) training procedure compared to $t$-SimCNE, directly learning the 2D embedding.
@@ 3.2 Trace uses subset means as positive pairs
We reasoned that the variation between responses in different trials provides an estimate for the naturally occurring variability at each time point, e.g. due to fluctuating brain state or inaccuracies in the measurement. This is precisely the type of noise to which the embedding should be invariant. Therefore, we use averages of two random, non-overlapping subsets of the trials as positive pairs, making the representation invariant to the trial-to-trial fluctuations.
Formally, let $r$ be the total number of trials, which are randomly split into two equisized, non-overlapping subsets $S_{1}$ and $S_{2}$ with
$S_{1}\cap S_{2}=\emptyset\quad\text{and}\quad|S_{1}|=|S_{2}|=k\leq r.$ (4)
Then the positive pair consists of
$y_{i}=\frac{1}{k}\sum_{l\in S_{1}}x_{i}^{l}\quad\text{and}\quad y^{\prime}_{i}=\frac{1}{k}\sum_{l\in S_{2}}x_{i}^{l}.$ (5)
This approach is inspired by mix-up-based augmentations [45], but generalized to averages of more than one sample and restricted to mixing only trial responses of the same neuron. The number $k$ of trials to average over is a hyperparameter, but we found that choosing $k\approx r/2$ leads to a good trade-off between variability of the positive pairs, and similarity to the full mean which we embed at inference time. It also leads to the largest number of distinct positive pairs, $\binom{r}{\lfloor r/2\rfloor}$. The subset means from other neurons in the batch were used as negative pairs.
Figure 3: Neural dataset from retinal ganglion-cells measured in superior colliculus and its embedding across different methods. (a) Schematic of recording neural activity. Each neuron projecting to superior colliculus (left) has many retinal boutons (middle) from which we record activity in response to a full-field light stimulus using two-photon calcium imaging (right). (b) We functionally clustered the light-evoked neural activity into four major groups (OFF, ON-OFF, ON, Suppressed-by-contrast, Sbc) showing their mean response to the “chirp” (left) and the ”moving bar” stimulus (right). (c) Trace embedding of the neuronal time series color-coded according to their functional group. (d) Same as (c), but for CEED visualized in 2D using PCA, (e) TS2Vec visualized in 2D using PCA, (f) $t$-SNE, and (g) UMAP.

### v2

@@ Abstract
Modern neural recording techniques such as two-photon imaging or Neuropixel probes allow to acquire vast time-series datasets with responses of hundreds or thousands of neurons. Contrastive learning is a powerful self-supervised framework for learning representations of complex datasets. Existing applications for neural time series rely on generic data augmentations and do not exploit the multi-trial data structure inherent in many neural datasets. Here we present Trace, a new contrastive learning framework that averages across different subsets of trials to generate positive pairs. Trace allows to directly learn a two-dimensional embedding, combining ideas from contrastive learning and neighbor embeddings. We show that Trace outperforms other methods, resolving fine response differences in simulated data. Further, using in vivo recordings, we show that the representations learned by Trace capture both biologically relevant continuous variation, cell-type-related cluster structure, and can assist data quality control.
@@ 1 Introduction
With advances in recording techniques, datasets in neuroscience have grown in size and complexity [chen2017neural, stringer2024analysis]. For example, two-photon imaging and Neuropixel probes have made it possible to record responses of tens of thousands of neurons from multiple cortical areas under comparable conditions [siegle2021survey, de2020large]. To summarize and visually explore such noisy, high-dimensional data, it is invaluable to represent it in two dimensions to identify functionally similar groups of neurons [zeng2017neuronal].
@@
A prominent paradigm for learning informative representations of data is contrastive learning [gutmann2012noise, oord2018representation]. Here, representations are created by contrasting similar samples (referred to as “positive pairs”) against dissimilar ones (“negative pairs”), ensuring that similar samples are grouped together, while dissimilar ones are separated. Contrastive learning has been popular for image data for some years [chen2020simple], but has only recently seen first applications for neuroscience time-series data [schneider2023learnable, vishnubhotla2023towards]. For example, the contrastive method CEED [vishnubhotla2023towards] learns representations of extracellular action potential wave forms using generic data augmentations such as amplitude jitter. However, most contrastive learning methods, including CEED, do not directly embed into two dimensions necessary for visualization.
@@
Here, we exploit a common feature of neuroscience experiments – multiple recorded responses to identical stimuli – to develop a new contrastive learning framework tailored for multi-trial time series in neuroscience. This framework directly learns a two-dimensional (2D) representation of the data. Building on $t$-SimCNE [bohmunsupervised], we present TraceTime series Representation Analysis through Contrastive Embeddings. Instead of using hand-tuned or generic data augmentations, Trace creates positive pairs by averaging across different subsets of trials, better capturing the structure of time-series variation in neural responses (Fig. 1).
@@
Figure 1: Trace embeds multi-trial time series using subset means for positive-pair generation. Left: Typical experimental structure in neuroscience: multiple trials of neural activity in response to repeats of an identical stimulus (here, full-field light intensity modulations). For Trace, positive pairs are generated using means of subsets of trials. Middle: Positive pairs are fed through an MLP and a fully-connected projection head to get representations $z^{\prime}_{i}$ and $z^{\prime\prime}_{i}$. The loss function pushes $z^{\prime}_{i}$ and $z^{\prime\prime}_{i}$ together and maximizes their similarity. Right: Final embedding of large-scale neuroscience dataset recorded in superior colliculus. Color-coded according to their functional group.
@@
Using a synthetic dataset, we show that Trace indeed automatically identifies informative parts of a time series, where other methods are “drowning in noise”. We then investigate the performance of Trace on a real-world two-photon-imaging and a Neuropixels time-series dataset. Trace produces visualizations that capture biological properties of the recorded cells better than other visualization techniques. Our code is available at https://github.com/berenslab/TRACE and https://github.com/berenslab/TRACE_experiments.
@@ 2 Related work
Towards this goal, two of the most prominent non-linear visualization methods are $t$-SNE [van2008visualizing] and UMAP [mcinnes2018umap]. While mostly used for single-cell transcriptomics data, they have also been employed for wave-shape classification for electrophysiological data [lee2021non]. $t$-SNE and UMAP learn a non-parametric 2D embedding, guided by pairs of nearest neighbors in data space. However, nearest neighbors in the high-dimensional space can be a poor proxy for semantic similarity due to the curse of dimensionality [hinneburg2000nearest]. Alternatively, self-supervised learning approaches train encoder networks to produce high-dimensional representations. For visual exploration, these need to be further reduced, e.g. using PCA [pearson1901liii]. Instead of pairs of nearest neighbors, these methods rely on data augmentations, better capturing the semantic similarities of samples. The type of augmentation depends on the data modality, e.g. randomly resized crops and flips for images [chen2020simple] or sampling consecutive time steps in speech recordings [oord2018representation]. Notably, $t$-SimCNE [bohmunsupervised] uses data augmentations to create positive pairs of images but embeds into 2D to visualize the data.
@@
Contrastive learning has been applied to medical time-series data, in particular electrocorticogram (ECoG) data. In addition to general time-series augmentations like scaling, blurring [sarkar2020self], and jittering [eldele2021time], frequency- [zhang2022self], cutout- [cheng2020subject], and permutation-based augmentations [sarkar2020self] have been used, which are not applicable to multi-trial non-periodic, stimulus-response data. Some of these works [kiyasseh2021clocs, diamant2022patient] also create positive pairs based on patient identity. Mix-up [zhang2018mixup] creates augmentations by blending different samples [bachman2019learning], which was also applied to ECoG data [cheng2020subject, wickstrom2022mixing]. Patient-based positive pairs or mix-up are similar in spirit to our approach. An important difference is that they only mix or pair two single trial samples and not a larger subset of trials, as we do. Our approach balances positive pair variability and similarity to inference-time inputs.
@@
In neuroscience, there has been comparatively little work on contrastive learning for time-series representations with the goal of identifying discrete cell types (i.e. clusters). peterson2022learning use a self-supervised model to generate pseudo labels and enrich the collected data this way. cho2023neural suggest a time-warping loss, which is not applicable to our experiments, which are aligned in time by construction. Related to our approach, the method CEED has been used for spike sorting and cell-type classification [vishnubhotla2023towards]. CEED generates representations using generic time-series as well as task-specific augmentations such as spike collision and channel subset selection. As is common for contrastive self-supervised learning, CEED does not embed into two dimensions directly, but relies on an additional dimensionality reduction step for a 2D visualization, unlike our method. Moreover, Trace captures the local noise structure better than CEED’s generic augmentations (Fig. 2).
@@
For learning representations of neural population dynamics, methods like LFADS [pandarinath2018inferring], SwapVAE [liu2021drop], or the contrastive approaches MYOW [azabou2021mine], CEBRA [schneider2023learnable], SinkDivLM [urzay2023detecting], and Neuroformer [antoniadesneuroformer] produce embeddings in which the temporal evolution of the population activity and behavioral modulation of neural activity can be explored. Here, each embedding point corresponds to a time point, not a single-neuron response like in Trace. Thus, they tackle a fundamentally different problem and are not applicable in our setting (see point (2) in the beginning of this section).
@@
Similarly, many general-purpose contrastive time-series models, like TNC [tonekaboniunsupervised], also produce time-point embeddings and are thus not applicable in our setting. Others, such as TS2Vec [yue2022ts2vec] contrast instances of time series, but learn high-dimensional embeddings, not geared towards visualization. We find that Trace outperforms TS2Vec even when adapting the latter to a visualization setting.
@@
NEMO [yuvivo] and PhysMAP [lee2024physmap] contrast different electorphysiological modalities, while Trace works by contrasting multiple recordings of the same modality.
@@ 3 Trace: contrastive learning for multi-trial time-series data in neuroscience
Our goal is to create informative 2D embeddings of multi-trial time-series data from neuroscience experiments, supporting data exploration, cell-type discovery, and other clustering tasks. To do so, we developed Trace, a method that makes use of the trial-based nature of many neuroscience experiments to create positive pairs by averaging over subsets of trials. It builds on the self-supervised method $t$-SimCNE [bohmunsupervised], which is based on the SimCLR framework [chen2020simple] and directly learns a 2D embedding. We briefly review contrastive learning and $t$-SimCNE before describing our approach. Finally, we will briefly describe the contrastive learning frameworks CEED and TS2Vec as our closest competitors using generic data augmentations for time series [vishnubhotla2023towards].
Figure 2: Synthetic dataset shows ability of Trace to identify cell types in noisy data. (a) Simulation of two classes of neural responses with baseline (25 s) and type-specific response (5 s) periods. In the baseline period, both classes show 10 Hz activity, while in the type-specific response period they fire at 11 Hz and 9 Hz, respectively. Responses are corrupted by variable amounts of noise (Sec. 4.1). (b) Discriminability [egan1962psychophysics] of the two neuron classes per bin was measured using $D=\frac{|\mu_{1}-\mu_{2}|}{0.5(\sigma_{1}+\sigma_{2})}$ (mean across neurons shown). (c) Mean neural responses across 10 trials for two example neurons from the two types. (d) Distributions of baseline (dashed) and response periods (solid) for the two neuronal responses in (c). (e – i) Embeddings of 10k simulated neurons for baseline activity noise SD={}={}38 with 10 trials each. (j) Discriminability for simulated responses with increasing amounts of noise during the neural baseline activity. Error bars indicate 95% confidence intervals.
@@ 3.1 Contrastive self-supervised visualizations
Given a dataset in a data space $X$, contrastive self-supervised learning trains the parameters $\theta$ of a neural network $f_{\theta}:X\to Z$ to obtain representations $z=f_{\theta}(x)$ in embedding space $Z$. It learns salient features by making the embedding invariant to known data similarities encoded in positive pairs, which are typically obtained as two modifications $^{\prime}, x′′x^{\prime\prime}$ of the same data sample $x$. Similarity in the embedding space is measured by a similarity function $q:Z\times Z\to\mathbb{R}$, Eqs. (2, 3). Positive pairs $(x^{\prime},x^{\prime\prime})$ should have high similarity $q(z^{\prime},z^{\prime\prime})$, i.e. $z^{\prime}$ and $z^{\prime\prime}$ should be close. This is typically achieved with the InfoNCE loss, which for the $i$-th positive pair $(x^{\prime}_{i},x^{\prime\prime}_{i})$ of the training batch is
$\mathcal{L}(x^{\prime}_{i},x^{\prime\prime}_{i})=-\log\frac{q(z^{\prime}_{i},z^{\prime\prime}_{i})}{q(z^{\prime}_{i},z^{\prime\prime}_{i})+\sum_{\alpha\neq i}\big(q(z^{\prime}_{i},z^{\prime}_{\alpha})+q(z^{\prime}_{i},z^{\prime\prime}_{\alpha})\big)}.$ (1)
Here, $\alpha$ runs over all other positive pairs in the training batch. The pairs $(z^{\prime}_{i},z^{\prime}_{\alpha})$ and $(z^{\prime}_{i},z^{\prime\prime}_{\alpha})$ are called negative pairs and their similarity is decreased, i.e. $z^{\prime}_{i}$ and $z^{\prime}_{\alpha}$ (or $z^{\prime\prime}_{\alpha}$) pushed apart.
Usually, the embedding space is constrained to a high-dimensional hypersphere $Z=\mathbb{S}^{d-1}\subset\mathbb{R}^{d}$ and the similarity function for $z,\tilde{z}\in Z$ is based on the cosine similarity:
$q(z,\tilde{z})=\exp\left((z^{\top}\tilde{z})/(\|z\|\|\tilde{z}\|\tau)\right),$ (2)
@@
$t$-SimCNE [bohmunsupervised] learns a 2D visualization of image datasets with a parametric encoder and data augmentations as similarity source. The key ingredient is the $t$-SNE-inspired Cauchy kernel
$q(z,\tilde{z})=(1+\|z-\tilde{z}\|^{2})^{-1}$ (3)
in the 2D embedding space $Z=\mathbb{R}^{2}$. We adapt $t$-SimCNE to multi-trial time-series data from neuroscience by implementing positive pairs based on trials, replacing the original image-based transformations (see below). We adopt a simpler (one-stage) training procedure compared to $t$-SimCNE, directly learning the 2D embedding.
@@ 3.2 Trace uses subset means as positive pairs
We reasoned that the variation between responses in different trials provides an estimate for the naturally occurring variability at each time point, e.g. due to fluctuating brain state or inaccuracies in the measurement. This is precisely the type of noise to which the embedding should be invariant. Therefore, we use averages of two random, non-overlapping subsets of the trials as positive pairs, making the representation invariant to the trial-to-trial fluctuations. Formally, let $r$ be the total number of trials, which are randomly split into two equisized, non-overlapping subsets $S_{1}$ and $S_{2}$ with
$S_{1}\cap S_{2}=\emptyset\quad\text{and}\quad|S_{1}|=|S_{2}|=k\leq r/2.$ (4)
Then the positive pair of neuron $i$ consists of the means of these non-overlapping trial subsets:
$x^{\prime}_{i}=\frac{1}{k}\sum_{l\in S_{1}}x_{i}^{l}\quad\text{and}\quad x^{\prime\prime}_{i}=\frac{1}{k}\sum_{l\in S_{2}}x_{i}^{l}.$ (5)
The subsets $S_{1}$, $S_{2}$ are dynamically resampled per neuron in each epoch. A training batch is made up of the positive pairs of that batch’s neurons. This approach is inspired by mix-up augmentations [zhang2018mixup], but generalized to averages of more than two samples and restricted to mixing only trial responses of the same neuron. The number $k$ of trials to average over is a hyperparameter, but in practice we use $k=\lfloor r/2\rfloor$ as it provides a good trade-off between variability of the positive pairs, and similarity to the full mean which we embed at inference time. This choice also leads to the largest number of distinct positive pairs, $\binom{r}{\lfloor r/2\rfloor}$. The subset means from other neurons in the batch were used as negative pairs.
Figure 3: Neural dataset from retinal ganglion cells measured in superior colliculus and its embedding across different methods. (a) Schematic of recording neural activity. Each neuron projecting to superior colliculus (left) has many retinal boutons (middle) from which we record activity in response to a full-field light stimulus using two-photon calcium imaging (right). (b) We functionally clustered the light-evoked neural activity into four major groups (OFF, ON-OFF, ON, Suppressed-by-contrast, Sbc) showing their mean response to the “chirp” (left) and the ”moving bar” stimulus (right). (c) Trace embedding of the neuronal time series colored by their functional group, (d, e) Same as (c), but for CEED / TS2Vec visualized in 2D using PCA, (f) $t$-SNE, and (g) UMAP.